### v1

# Adaptive Alpha Weighting with PPO: Enhancing Prompt-Based LLM-Generated Alphas in Quant Trading

Abstract
This paper proposes a reinforcement learning framework that employs Proximal Policy Optimization (PPO) to dynamically optimize the weights of multiple large language model (LLM)-generated formulaic alphas for stock trading strategies. Formulaic alphas are mathematically defined trading signals derived from price, volume, sentiment, and other data. Although recent studies have shown that LLMs can generate diverse and effective alphas, a critical challenge lies in how to adaptively integrate them under varying market conditions. To address this gap, we leverage the deepseek-r1-distill-llama-70b model to generate fifty alphas for five major stocks: Apple, HSBC, Pepsi, Toyota, and Tencent, and then use PPO to adjust their weights in real time. Experimental results demonstrate that the PPO-optimized strategy achieves strong returns and high Sharpe ratios across most stocks, outperforming both an equal-weighted alpha portfolio and traditional benchmarks such as the Nikkei 225, S&P 500, and Hang Seng Index. The findings highlight the importance of reinforcement learning in the allocation of alpha weights and show the potential of combining LLM-generated signals with adaptive optimization for robust financial forecasting and trading.
Keywords Formulaic Alpha Generation  $\cdot$ LLM  $\cdot$ Proximal Policy Optimization  $\cdot$ Stock Prediction  $\cdot$ Time Series Forecasting
1 Introduction
Over the past few decades, financial markets have evolved significantly with the integration of advanced technologies, especially in quantitative finance. Traditionally, stock trading strategies have relied on financial indicators such as moving averages, volatility measures, and momentum indicators to predict market behavior and generate profits 9580172 ; MOSTAFAVI2025100631 . These methods have been successful to some extent, but often lack the flexibility and adaptability required to respond to changing market conditions. In addition, relying solely on manually chosen indicators may restrict the discovery of new patterns within the extensive and constantly changing financial data landscape.
Furthermore, trading signals often lose their effectiveness over time due to changes in market conditions and investor behavior 8279188 . This phenomenon is called Alpha Decay in quantitative finance. As a result, traders and investors must continuously search for new signals or develop new features that remain closely related to stock price movements. Staying adaptive and proactive is essential to maintain an edge in the dynamic and highly competitive financial markets.
To deal with rapidly changing and complex financial markets, advanced models and techniques have been developed to manage the increasing volume and diversity of data. One of the most promising advances in recent years is the use of machine learning, particularly deep learning models, in the development of trading strategies MINTARYA202396 . These models can process large datasets and uncover complex patterns in market behavior that traditional methods might miss. However, despite their impressive capabilities, machine learning models still face challenges in areas such as interpretability Sejnowski_2020 , flexibility, and the need for extensive data labeling 10114634 ; kou2024automatestrategyfindingllm .
A new development in this field involves the application of large language models (LLMs), such as OpenAI’s GPT models, for financial analysis. LLMs, which have been trained on large amounts of text data, can process financial news, reports, and historical price information to generate insights and predictions 10825946 . These models can generate a variety of outputs, including summarizing financial news and creating algorithmic strategies for trading. In particular, prompt-based LLMs, which take in specific prompts and generate relevant outputs based on those inputs, provide a powerful tool for automating and enhancing the alpha generation process in stock trading wang2024gptsignalgenerativeaisemiautomated . However, most of the time, other researchers only use LLMs to generate formulaic alphas and examine their correlations with stock return. They do not analyze how the LLM-generated alphas can be used in dynamic quant trading, which will be further investigated in this paper.
In our previous study chen2025sentimentawarestockpriceprediction , LLM-generated formulaic alphas have been demonstrated to have strong predictive power in stock price prediction. In this study, we extend this line of research by shifting the focus from alpha generation to alpha integration and optimization. Instead of relying solely on the predictive strength of individual alphas, we investigate how reinforcement learning, specifically Proximal Policy Optimization, can be employed to dynamically allocate weights across multiple LLM-generated alphas. This approach enables the trading strategy to adapt to evolving market conditions, enhance robustness against noisy signals, and capture diverse market dynamics.
First, we use a prompt-based LLM, deepseek-r1-distill-llama-70b model, to generate formulaic alphas for five major companies: Toyota, Apple, HSBC, Pepsi and Tencent. Alpha, in strict financial definition, refers to the excess return that a trading strategy can achieve above a benchmark index, such as the S&P 500 or Nikkei 225 kakushadze2016101formulaicalphas . In quantitative finance and machine learning, alpha factors often refer to signals or features that are predictive of future stock returns. Formulaic alpha involves mathematically defined trading signals or strategies aimed at predicting future asset returns. These strategies are usually expressed as explicit formulas that combine various types of data, including market, fundamental, or alternative inputs such as price, volume, volatility, and sentiment indicators. Traditionally, these formulas often originate from quantitative research. Equation 1 is an example of formulaic alpha. When a stock price is substantially higher than its recent average, equation 1 yields a negative value, indicating that the stock might be overbought and likely to decline. However, if the price is below the average, the alpha becomes positive, suggesting a potential buying opportunity due to mean reversion.
$\alpha_{t}=-(\text{Close}_{t}-\text{SMA}_{10,t})$ (1)
  • $\text{Close}_{t}$: The closing price of the stock at time $t$.
  • $\text{SMA}_{10,t}$: The 10-day Simple Moving Average of the stock’s closing price at time $t$.
Once formulaic alphas are generated, this study applies proximal policy optimization (PPO) to optimize the weights of each alpha. The reason for combining LLM-generated alphas with reinforcement learning is that recent studies often treat each LLM-generated alpha independently or combine them using static or manually designed weighting schemes. However, financial markets are dynamic and often non-stationary, meaning that the effectiveness of individual alphas can vary significantly over time. To address this, we propose combining LLM-generated alphas with reinforcement learning to dynamically learn optimal alpha weights based on changing market conditions. This integration allows the model to treat alpha weighting as a sequential decision-making task, where the agent continuously adjusts its strategy in response to observed market feedback. Compared to traditional methods such as regression-based weighting or simple averaging, our approach enables the system to adaptively prioritize more informative or better performing alphas in real time, thus improving overall trading performance and robustness.
PPO, a model-free algorithm proposed by schulman2017proximalpolicyoptimizationalgorithms , is particularly well-suited for continuous action spaces, making it ideal for determining the optimal weight distribution of different alphas in a portfolio. By training the PPO agent, this study seeks to identify the best combination of alphas that can maximize returns.
To evaluate the performance of our strategy, we compare the returns generated by the optimized alphas with those from an equal-weighted portfolio and major financial benchmarks, including the Nikkei 225, S&P 500, and Hang Seng Index. For example, the Nikkei 225 is one of the main stock market indices in Japan, tracking 225 large publicly traded companies in diverse sectors such as technology, automotive, finance and consumer goods. These benchmarks are widely used to represent the broader market and serve as reference points for assessing the effectiveness of trading strategies. In addition to cumulative returns, we also analyze the Sharpe ratio and maximum drawdown to provide a more comprehensive view of performance. The Sharpe ratio reflects the risk-adjusted return of the strategy, helping to assess whether higher returns are achieved efficiently. Maximum drawdown captures the largest peak-to-trough decline, indicating potential risk exposure during unfavorable market conditions. By evaluating these metrics together, we aim to determine not only whether LLM-generated alphas can outperform traditional market indices but also whether they can do so with more favorable risk-return characteristics.
The primary goal of this research is to demonstrate the potential of combining LLMs for alpha generation with reinforcement learning techniques such as PPO to create more adaptive and dynamic trading strategies. By exploring how machine learning models can work together with financial theory, this study offers insights into how technology can transform financial decision-making and trading practices.
The main contributions of this paper are as follows.
  1. 1.
    This study introduces a reinforcement learning framework using PPO to dynamically optimize the weights of multiple LLM-generated alphas, adapting to changing market conditions.
  2. 2.
    Empirical analysis is conducted across multiple alpha selection strategies, including random, low-correlation, and high-contribution methods, demonstrating that the proposed framework consistently outperforms market benchmarks across different stocks. This provides practical insights into how alpha quality and selection criteria influence portfolio performance.
The remainder of this paper is structured as follows. Firstly, related work is listed (Section 2). The methodology is then described (Section 3). In Section 4, the experimental results are presented. Further discussion is provided in Section 5. Finally, Section 6 concludes the paper.
2 Related Work
The concept of alpha generation in finance refers to the ability of a trading strategy to outperform the market or a benchmark, such as stock indices. Traditional alpha generation methods are based on fundamental and technical analysis, which identified patterns in historical stock prices, trading volumes, and other economic factors. However, as financial markets become more complex, quantitative approaches leveraging machine learning, and more recently deep learning, have been increasingly used to develop formulaic alphas, which are mathematical formulas or models designed to predict market movements and generate excess returns.
Early efforts in formulaic alpha generation rely on statistical and econometric models. For example, Fama and French FAMA19933 propose factor models, which aim to explain asset returns using factors such as market risk, size, value, and momentum. These models have been foundational in understanding systematic risk factors, but they have limitations in capturing complex, non-linear relationships in financial data. More recent advancements have focused on enhancing traditional alpha models by incorporating machine learning algorithms.
Machine learning approaches have gained significant attention in the domain of alpha generation. One notable development is the use of supervised learning models to predict stock prices and returns. Researchers have applied various machine learning techniques, including decision trees 8250694 , support vector machines (SVMs) 6703096 ; 9361804 , and random forests 9987903 , to identify patterns in historical data and generate alpha signals. However, these models are often limited by their reliance on structured data and the need for manual feature engineering, which can be time-consuming and prone to biases.
A more recent shift in alpha generation research has involved the application of deep learning techniques, particularly convolutional neural networks (CNNs) 9317207 and recurrent neural networks (RNNs) 10392023 , to stock market prediction. These models are capable of automatically learning complex and non-linear patterns from raw market data. Recently, some researchers have proposed some hybrid models to further improve prediction performance. For example, Zhang et al. math11091985 propose a CNN-BiLSTM-Attention-based model to improve the performance of stock price forecasting. The proposed approach can outperform models such as LSTM. Lu et al. https://doi.org/10.1155/2020/6622927 propose a CNN-LSTM to predict the stock price.
Recently, transformer-based models have become popular candidates in stock price forecasting. Some transformer models designed for time series forecasting include Dual Transformer chen2025sentiment , Informer zhou2021informerefficienttransformerlong , Autoformer wu2022autoformerdecompositiontransformersautocorrelation and TimeXer wang2024timexerempoweringtransformerstime . Li et al. li2023mastermarketguidedstocktransformer proposed a market-guided transformer to predict the stock price. This model uses momentary and cross-time stock correlation as well as market information for automatic feature selection.
In addition to supervised learning and deep learning, natural language processing (NLP) techniques have been used to extract information from unstructured data sources such as financial news articles, earnings reports, and social media. NLP models, especially generative LLMs such as LLaMA, Mistral, and GPT, have shown promise in sentiment analysis and in predicting stock movements based on textual data 10825946 ; yang2023fingptopensourcefinanciallarge ; xie2023pixiulargelanguagemodel . These models analyze the sentiment of news and social media content and generate alpha signals based on the prevailing market sentiment. This approach has become more prominent due to the increasing availability of financial news and the growing importance of market sentiment in stock price movements.
Prompt-based LLMs are another emerging area of interest in formulaic alpha generation. These models generate specific financial alphas based on customized input prompts. For example, researchers have explored the use of prompt-based LLMs to generate alphas that incorporate technical indicators, such as moving averages, RSI, and MACD, along with sentiment data from news articles or earnings calls kou2025automatestrategyfindingllm . For example, Chen and Kawashima chen2025sentimentawarestockpriceprediction utilize an LLM to generate formulaic alphas and provide evidence of their effectiveness in predicting stock price movements. Wang et al. wang2024gptsignalgenerativeaisemiautomated use the prompt-based GPT-4 model to generate formulaic alphas. By tailoring the prompts to include specific financial metrics, these LLMs can generate alphas that are directly aligned with the desired prediction task. This approach differs from traditional machine learning techniques because it combines domain-specific knowledge (in the form of technical indicators) with the generative capabilities of LLMs to create actionable trading strategies. However, their research only generates alphas but does not address how to adaptively combine them or assess their performance over time.
Reinforcement learning (RL) has gained increasing attention in financial forecasting and trading. RL models are designed to learn optimal trading strategies by interacting with the market environment through trial and error. By continuously updating their policies based on observed rewards, RL algorithms can adapt to changing market dynamics and improve decision making over time Chen2025 . Various RL methods, such as Deep Q-Networks (DQN), Proximal Policy Optimization (PPO), and Soft Actor-Critic (SAC), have been successfully applied to tasks like portfolio management, asset allocation, and algorithmic trading huang2025deepreinforcementlearningframework ; ndikum2024advancinginvestmentfrontiersindustrygrade ; 10627674 ; ijfs11010010 . These approaches allow trading agents to make sequential decisions, balancing the trade-off between risk and return in complex and uncertain markets.
The integration of LLMs and RL for optimizing multiple formulaic alphas is still a relatively novel area of research. Few studies have directly addressed the use of LLMs for alpha generation in combination with RL algorithms for weight optimization. We will examine the effectiveness of this framework in this study.
3 Methodology
In this study, we focus on generating formulaic alphas for predicting stock returns using a combination of prompt-based LLMs and reinforcement learning to optimize the weights of each alpha. The goal is to develop a framework that can produce alpha signals for multiple companies and optimize these signals to improve trading performance. Figure 2 illustrates the setup of this research.
Figure 1: Flow Chart of the Proposed Method
Figure 2: Whole Picture of the Proposed Method
3.1 Data
The data used for this research come from two primary channels: stock data and financial news sentiment data.
Historical stock price data for Toyota, Apple, HSBC, Pepsi, and Tencent from 2016/02/16 to 2024/05/08 are downloaded using the yfinance Python library. This library provides an efficient way to access daily stock prices, including opening, closing, high, low, and volume data for each company. Stock data is used to calculate key technical indicators, such as moving averages, momentum, and RSI, which are incorporated into the alpha generation process. The Python library pandas-ta is used to calculate all the technical indicators. In finance, technical indicators, calculated from historical stock price and volume data, are often used by traders and investors to predict the stock trend and generate trading signals. Table 1 shows the technical indicators used in this study.
The Simple Moving Average (SMA) calculates the average price of a security over a specified period, smoothing out short-term fluctuations to reveal overall trends. In contrast, the Exponential Moving Average (EMA) places greater emphasis on recent prices, making it more responsive to new information compared to the SMA. Momentum measures the speed of price changes by comparing the current price to a previous price from a set number of periods ago, helping traders identify potential trend changes. The Relative Strength Index (RSI) is a momentum oscillator that assesses the speed and change of price movements, commonly used to detect overbought or oversold conditions. MACD (Moving Average Convergence Divergence) tracks the relationship between two EMAs, helping traders understand changes in the strength, direction, and momentum of a trend. Bollinger Bands are plotted two standard deviations away from a moving average, offering a visual representation of price volatility and potential reversal points. On-Balance Volume (OBV) combines price movement with volume flow to indicate how volume affects the direction of the price.
Figure 3: Close Prices of All Five Companies
Table 1: Technical Indicators Used for Alpha Generation
Indicator Name
Simple Moving Average (5-day)
Simple Moving Average (20-day)
Exponential Moving Average (10-day)
Momentum (3-day)
Momentum (10-day)
Relative Strength Index (14-day)
Moving Average Convergence Divergence (MACD)
MACD Signal Line
Bollinger Band Upper
Bollinger Band Lower
On-Balance Volume (OBV)
Daily news articles related to target companies are downloaded from websites such as Yahoo News using the Financial News Feed and Stock News Sentiment Data API  [^1] https://eodhd.com/financial-apis/stock-market-financial-news-api. The sentiment of these articles is computed using a natural language processing model, NLTK that analyzes the tone of the text and outputs a polarity score. The polarity score is used to quantify sentiment, ranging from 1-1 to $1$. A score close to $1$ indicates strong positive sentiment, while a score close to 1-1 reflects strong negative sentiment. A score of 0 represents neutral sentiment. This sentiment score is then used to influence the generation of alphas, as market sentiment can significantly impact stock prices. The sentiment data are integrated into the LLM prompt, along with historical price and technical indicator data, to generate formulaic alphas for each company.
The first 80% of the data will be used for training and the remaining 20% of the data will be used for testing.
3.2 Model Setup
3.2.1 Prompt-Based LLM for Alpha Generation
The model used to generate formulaic alphas is a prompt-based LLM, specifically deepseek-r1-distill-llama-70b. This LLM is from the American AI company Groq [^2] https://groq.com/. This version of DeepSeek model has been shown to generate competitive results without fine-tuning compared to other LLMs such as LLaMA and Gemma in previous study Chen2025 . This model (deepseek-r1-distill-llama-70b) uses knowledge distillation, a process in which a smaller model (the student) learns from the output of a larger, more complex model (the teacher). In this case, the LLaMA-70B model serves as the teacher, and the student model is trained to replicate its behavior. The teacher model generates predictions or soft targets, which guide the student in learning more efficiently. Knowledge distillation allows the student model to retain much of the performance of the teacher model while being more lightweight, making it suitable for deployment in environments with limited resources hinton2015distillingknowledgeneuralnetwork .
The input data fed to the prompt ({features} in Table 2) include historical prices, technical indicators, and sentiment scores of all companies. The input data are stored in a pandas DataFrame and then converted to JSON format. The model generates a common set of fifty distinct formulaic alphas shared across all companies. These alphas are mathematical expressions that combine various financial metrics, such as momentum, moving averages, RSI, MACD, and Bollinger Bands. The LLM is prompted with historical data and a detailed description of the task, which guide it to generate alphas that can be used to predict stock movements.
For example, one formula might combine momentum indicators with moving averages, while another might include sentiment scores to account for market news. Fifty formulas are generated for the forecasting of stock returns of all companies.
Table 2: LLM Prompt for Formulaic Alpha Generation
You are a quantitative trader. Generate 50 alpha formulas using the given stock features:
{features}.
The formulas should be mathematical expressions combining these features.
Return only the formulas in Python syntax, using variables like C_t (Close), O_t (Open),
V_t (Volume), S_t (Sentiment), and standard indicators (SMA, Momentum).
Example Output: alpha_t = (C_t - O_t) / O_t + 0.5 * S_t
3.2.2 PPO for Alpha Weight Optimization
The alpha weight optimization process employs PPO, a reinforcement learning algorithm particularly suited for continuous action spaces in financial applications. This choice is motivated by PPO’s three key properties: (1) its clipped objective function prevents destructively large policy updates, (2) its sample efficiency aligns with the limited availability of high-quality financial data, and (3) its ability to handle the non-stationary dynamics inherent in market environments.
The trading environment constitutes a partially observable Markov decision process (POMDP) where the state space $s_{t}\in\mathcal{S}$ at time $t$ integrates four information categories:
$s_{t}=\{\text{OHLCV}_{t},\,p_{t-1},\,\text{regime}_{t},\,\sigma_{t}\}$ (2)
Here, $\text{OHLCV}_{t}$ (Open, High, Low, Close, Volume) represents the raw price and volume data, $p_{t-1}\in[-1,1]$ denotes the previous position, $\text{regime}_{t}\in\{0,1\}$ indicates the bull/bear market classification based on the 20-day and 100-day moving average crossover, and $\sigma_{t}$ captures the 63-day rolling daily volatility annualized via $\sigma_{t}^{\text{annual}}=\sigma_{t}^{\text{63d}}\times\sqrt{252}$. The action space $\mathcal{A}$ consists of 50-dimensional weight vectors $\mathbf{w}_{t}\in\mathbb{R}^{50}$ corresponding to each alpha signal. These weights undergo normalization through an $L_{1}$-constrained transformation:
$\tilde{\mathbf{w}}_{t}=\operatorname{clip}(\mathbf{w}_{t},-1,1)$ (3)
$\mathbf{w}_{t}^{\text{norm}}=\frac{\tilde{\mathbf{w}}_{t}}{\|\tilde{\mathbf{w}}_{t}\|_{1}+\epsilon},\qquad\epsilon=10^{-8}$ (4)
Clipping ensures that each element of the weight vector $\mathbf{w}_{t}$ lie in the interval [1,1][-1,1], which prevents extreme weights and enforces a bounded action space at the individual weight level. After clipping, the weight vector is divided by its $L_{1}$-norm $\|\tilde{\mathbf{w}}_{t}\|_{1}$, which ensures that the sum of the absolute values of the normalized weights is approximately 1. This is like requiring the agent to allocate weights across signals such that the total exposure is normalized to 1. This normalization preserves the interpretability of long/short signals while maintaining numerical stability by adding the $\epsilon$ to avoid division by zero.
The reward function $r_{t}$ combines two critical components:
$r_{t}=\underbrace{p_{t}\cdot R_{t}^{\text{future}}}_{\text{Position P\&L}}-\underbrace{\lambda|p_{t}-p_{t-1}|}_{\text{Transaction Cost}}$ (5)
where $R_{t}^{\text{future}}$ denotes the realized return of the asset over the next period, measured as the relative price change from $t$ to $t+1$. $\lambda=0.1\%$ quantifies the transaction cost. The position $p_{t}$ derives from a multi-stage calculation
$\alpha^{\text{composite}}_{t}=\sum_{i=1}^{50}w_{t}[i]\cdot\alpha_{i,t}$ (6)
$p_{t}=v_{t}\cdot\begin{cases}\min(1,2(\alpha_{t}^{\text{composite}}-\tau_{t}^{\text{upper}}))&\text{if }\alpha_{t}^{\text{composite}}>\tau_{t}^{\text{upper}}\\ \max(-1,2(\alpha_{t}^{\text{composite}}-\tau_{t}^{\text{lower}}))&\text{if }\alpha_{t}^{\text{composite}}<\tau_{t}^{\text{lower}}\\ 0&\text{otherwise}\end{cases}$ (7)
with $v_{t}=\min(2,\sigma_{\text{target}}/\sigma_{t}^{\text{annual}})$ implementing volatility scaling ($\sigma_{\text{target}}=0.15$) and $\tau_{t}$ representing quintile thresholds. $\alpha_{t}^{\text{composite}}$ is the composite alpha, computed as a weighted sum of the individual alpha signals at time $t$ (Equation 6). Each individual alpha is standardized with the StandardScaler prior to computing the composite alpha.
The volatility scaling factor $v_{t}$ dynamically adjusts position sizes to maintain stable risk exposure. When realized volatility ($\sigma_{t}^{\text{annual}}$) exceeds the target 15%, positions are scaled down to prevent overexposure during turbulent markets. Conversely, positions can scale up in low-volatility regimes to capitalize on opportunities without exceeding the risk budget.
The rolling quintile thresholds $\tau_{t}^{\text{upper}}=Q_{0.75}(\text{Close}_{t-126:t})$ and $\tau_{t}^{\text{lower}}=Q_{0.25}(\text{Close}_{t-126:t})$ serve as adaptive signal filters. Using the 75th and 25th percentiles of recent prices, these thresholds automatically adjust to market conditions. They widen when prices swing wildly (needing stronger signals before trading) and tighten when prices move steadily (capturing smaller opportunities). This prevents excessive trading during unpredictable back-and-forth price movements while still responding to real trends. The 126-day window balances long enough to filter out temporary market noise but short enough to adapt to changing conditions.
The PPO objective function incorporates these elements through
$\mathcal{L}^{\text{CLIP}}(\theta)=\mathbb{E}_{t}\left[\min\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\hat{A}_{t},\ \text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},\ 1-\epsilon,\ 1+\epsilon\right)\hat{A}_{t}\right)\right]$ (8)
where $\epsilon=0.2$ defines the clipping range and $\hat{A}_{t}$ denotes the advantage estimate computed using generalized advantage estimation (GAE). This formulation enables stable weight updates while accounting for the non-IID nature of financial time series through proper advantage discounting.
The PPO agent was trained using the default hyperparameters of Stable-Baselines3  [^3] https://stable-baselines3.readthedocs.io/en/master/, with learning rate $3\times 10^{-4}$, rollout length 2048, batch size 64, 10 update epochs per rollout, discount factor 0.99, and clipping parameter 0.2.
Table 3 shows the whole picture of the alpha trading algorithm and Table 4 illustrates the key parameters and their ranges of values.
Table 3: Alpha Trading Algorithm
Algorithm: Alpha Trading with Adaptive Risk Controls
Inputs:
 Price data $\mathcal{D}=\{OHLCV_{t},\text{FutureReturn}_{t}\}_{t=1}^{T}$
 Alpha signals $\{\alpha_{i,t}\}_{i=1}^{50}$
 Volatility target $\sigma_{\text{target}}=0.15$
 Transaction cost $\lambda=0.001$
Outputs:
 Performance metrics $\mathcal{M}$
Initialization:
1. Preprocess data:
$\alpha_{i}\leftarrow\text{StandardScaler.fit\_transform}(\alpha_{i}),\ \forall i\in 1:50$
$\text{MA}_{20}\leftarrow\text{rolling\_mean}(\text{Close},20)$
$\text{MA}_{100}\leftarrow\text{rolling\_mean}(\text{Close},100)$
$\text{regime}\leftarrow\mathbb{I}(\text{MA}_{20}>\text{MA}_{100})$
$\sigma_{\text{daily}}\leftarrow\text{rolling\_std}(\text{FutureReturn},63)$
$\sigma_{\text{annual}}\leftarrow\sigma_{\text{daily}}\times\sqrt{252}$
$\tau^{\text{upper}}\leftarrow\text{rolling\_quantile}(\text{Close},0.75,126)$
$\tau^{\text{lower}}\leftarrow\text{rolling\_quantile}(\text{Close},0.25,126)$
2. Initialize tracking: $V\leftarrow 1.0,\ p\leftarrow 0,\ \text{peak}\leftarrow 1.0$
Trading Loop:
3. for $t=1$ to T1T-1 do:
 4. Agent Decision:
  $\mathbf{w}_{t}\leftarrow\text{model.predict}(s_{t})$
  $\tilde{\mathbf{w}}_{t}\leftarrow\operatorname{clip}(\mathbf{w}_{t},-1,1)$
  $\mathbf{w}_{t}^{\text{norm}}\leftarrow\frac{\tilde{\mathbf{w}}_{t}}{\|\tilde{\mathbf{w}}_{t}\|_{1}+\epsilon},\quad\epsilon=10^{-8}$
 5. Composite Signal:
  $\alpha^{\text{composite}}_{t}\leftarrow\sum_{i=1}^{50}w_{t}[i]\cdot\alpha_{i,t}$
 6. Position Sizing:
  $p_{t}\leftarrow\begin{cases}\min(1,2(\alpha^{\text{composite}}_{t}-\tau^{\text{upper}}_{t})),&\alpha^{\text{composite}}_{t}>\tau^{\text{upper}}_{t}\\ \max(-1,2(\alpha^{\text{composite}}_{t}-\tau^{\text{lower}}_{t})),&\alpha^{\text{composite}}_{t}<\tau^{\text{lower}}_{t}\\ 0,&\text{otherwise}\end{cases}$
[…]
  Vol scaling: $v_{t}\leftarrow\min(2,\sigma_{\text{target}}/\sigma^{\text{annual}}_{t})$
  $p_{t}\leftarrow p_{t}\cdot v_{t}$
 7. Execute Trade:
  $\text{cost}\leftarrow\lambda\cdot|p_{t}-p_{t-1}|$
  $r_{t}\leftarrow p_{t}\cdot\text{FutureReturn}_{t}-\text{cost}$
  $V_{t}\leftarrow V_{t-1}\cdot(1+r_{t})$
  $\text{peak}\leftarrow\max(\text{peak},V_{t})$
  $\text{DD}_{t}\leftarrow(V_{t}-\text{peak})/\text{peak}$
 8. Update State:
  $s_{t+1}\leftarrow\{\text{OHLCV}_{t+1},p_{t},\text{regime}_{t+1},\sigma^{\text{daily}}_{t+1}\}$
end for
Performance Evaluation:
9. Compute metrics:
$\text{Sharpe}\leftarrow\frac{\mathbb{E}[R]}{\sigma_{R}}\times\sqrt{252}$
$\text{MaxDD}\leftarrow\min(\text{Cumulative}/\text{cumulative\_max}(\text{Cumulative})-1)$
Table 4: Key Parameters and Their Value Ranges
Parameter Description Typical Range / Value
$\sigma_{\text{target}}$ Annualized volatility target 0.15 (fixed)
$\lambda$ Transaction cost per trade 0.001
$\alpha_{i,t}$ Alpha signals Normalized to $\sim\mathcal{N}(0,1)$
$\epsilon$ Stability constant in normalization $10^{-8}$
$p_{t}$ Position sizing value [1,1][-1,1]
$v_{t}$ Volatility scaling factor $[0,2]$
$\tau^{\text{upper}}_{t}$ Upper threshold (75% quantile) Time-varying
$\tau^{\text{lower}}_{t}$ Lower threshold (25% quantile) Time-varying
$w_{t}[i]$ Alpha weight from model [1,1][-1,1], normalized
$\text{MA}_{20},\text{MA}_{100}$ Moving averages for regime filter Computed over 20 and 100 days
$\sigma_{\text{daily}}$ Daily return volatility Rolling 63-day std. dev.
$\sigma_{\text{annual}}$ Annualized volatility $\sigma_{\text{daily}}\times\sqrt{252}$
3.3 Evaluation
To evaluate the effectiveness of the proposed framework, we compare its trading performance with an equal-weighted strategy and major market indices. In the equal-weighted strategy, each LLM-generated alpha is assigned the same weight within the portfolio. The framework is considered effective if it consistently outperforms both the equal-weighted strategy and the market indices. To assess performance comprehensively, we use a range of metrics, including Information Coefficient (IC), Mutual Information, LightGBM Feature Importance, Cumulative Return, Maximum Drawdown, and Sharpe Ratio.
First, the Information Coefficient (IC) (Equation 9) is calculated for each alpha signal across the five selected companies: Apple, HSBC, Pepsi, Tencent, and Toyota. The IC measures the correlation between the predicted signals and the actual future returns. A positive IC value close to or above 0.05 indicates that the alpha has significant predictive power. IC values below 0.05 suggest weak predictive power, while negative IC values imply an inverse relationship, which may still be useful to construct contrarian strategies.
A contrarian strategy is an investment approach that opposes prevailing market trends. Contrarian investors buy when others are selling and sell when others are buying, anticipating price reversals dreman1998contrarian . In this context, a negative IC indicates that the signal moves in the opposite direction of return. Although the signal does not directly predict returns, it can still be valuable for contrarian strategies, where trading against the signal’s direction is intentional.
$IC=\frac{\text{Cov}(\text{rank}(\hat{r}),\text{rank}(r))}{\sigma_{\text{rank}(\hat{r})}\cdot\sigma_{\text{rank}(r)}}$ (9)
where:
  • $\hat{r}$ is the vector of predicted returns (alpha signals)
  • $r$ is the vector of realized future returns
  • $\text{rank}(\cdot)$ denotes the rank transformation (for Spearman correlation)
  • $\text{Cov}(\cdot,\cdot)$ is the covariance between ranked variables
  • $\sigma$ is the standard deviation of the ranked values
Mutual Information (MI) (Equation 10) quantifies the amount of information shared between an input feature $X$ and the target variable $Y$. It measures how much knowing $X$ reduces the uncertainty about $Y$, capturing both linear and nonlinear relationships. Mathematically, the mutual information is defined as:
$I(X;Y)=\sum_{x\in X}\sum_{y\in Y}p(x,y)\log\left(\frac{p(x,y)}{p(x)p(y)}\right)$ (10)
where $p(x,y)$ is the joint probability distribution of $X$ and $Y$, and $p(x)$, $p(y)$ are their marginal distributions. A higher MI value indicates a stronger dependence between the feature and the target. Features with near-zero MI contribute little to predicting output and can often be discarded.
LightGBM Feature Importance (Equation 11) is derived from the internal structure of the gradient boosting decision tree model. During training, LightGBM builds an ensemble of decision trees by sequentially minimizing a specified loss function. At each decision node, the model selects the feature that provides the greatest reduction in loss, commonly measured by information gain. The importance of a feature is then computed by summing its contribution to the loss reduction across all trees and relevant splits. This is formalized as follows.
$\text{Importance}(f_{i})=\sum_{t=1}^{T}\sum_{s\in\mathcal{S}_{t}(f_{i})}\Delta\mathcal{L}_{t,s}$ (11)
where $T$ is the total number of trees, $\mathcal{S}_{t}(f_{i})$ denotes the set of splits in tree $t$ that involve feature $f_{i}$, and $\Delta\mathcal{L}_{t,s}$ represents the reduction in the loss at split $s$.
LightGBM provides several types of feature importance metrics:
  • Gain Importance: Measures the total loss reduction brought by each feature across all splits. It reflects how much each feature contributes to the model’s performance and is considered the most informative metric.
  • Split Importance: Counts the number of times a feature is used in splits, regardless of the magnitude of improvement. While it captures feature frequency, it may not accurately reflect predictive value.
  • Cover Importance: Measures the number of samples affected by splits involving each feature. It helps assess how broadly a feature influences the dataset.
In this study, only the Gain Importance is reported and visualized for the LLM-generated alphas. This choice is motivated by the fact that gain importance directly quantifies a feature’s contribution to reducing prediction error. Since the primary objective is to identify which alphas most significantly improve model performance, gain importance offers a more relevant and interpretable measure. Other metrics, such as split and cover, might be influenced by frequent but weak splits or uneven sample distributions, making them less suitable for evaluating predictive strength.
The Cumulative Return of each company’s portfolio, optimized by the Proximal Policy Optimization (PPO) algorithm, is compared to its respective market benchmark (e.g., S&P 500 for Apple and Pepsi), Hang Seng Index for HSBC and Tencent, and Nikkei 225 for Toyota. This comparison assesses whether the RL-optimized alpha weights can outperform traditional market indices over the evaluation period.
$R_{\text{cum}}=\prod_{t=1}^{T}(1+r_{t})-1$ (12)
where:
  • $R_{\text{cum}}$ is the cumulative return over the time horizon $T$
  • $r_{t}$ is the return at time $t$
  • $\prod$ denotes the product over all time periods from $t=1$ to $T$
In addition to cumulative return, we employ the Sharpe Ratio to evaluate the risk-adjusted performance of our PPO-adjusted trading strategy. The Sharpe Ratio measures the excess return per unit of risk and is defined as:
$\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}$ (13)
where $R_{p}$ is the average return of the stock, $R_{f}$ is the risk-free rate (typically the return of a short-term government bond), and $\sigma_{p}$ is the standard deviation of stock returns. A higher Sharpe Ratio indicates that the strategy achieves better returns for each unit of risk taken. A Sharpe ratio higher than one is generally considered a sign of a good trading strategy, indicating that the strategy delivers returns significantly higher than the risk taken. This metric is particularly important when comparing strategies with different volatility profiles.
We also consider Maximum Drawdown (MDD) to evaluate downside risk. MDD represents the largest peak-to-trough decline in the value of a stock before a new peak is achieved. Formally, it is defined as:
$\text{MDD}=\max_{t\in[0,T]}\left(\frac{\max_{\tau\in[0,t]}P(\tau)-P(t)}{\max_{\tau\in[0,t]}P(\tau)}\right)$ (14)
where $P(t)$ denotes the stock value at time $t$. Max Drawdown captures the worst historical loss from a stock high and is critical for understanding the potential downside exposure of a strategy. Lower values of MDD are preferable, indicating better capital preservation under adverse market conditions.
The combined use of these evaluation metrics provides a comprehensive view of both the predictive strength of the alphas and the practical profitability of the final strategy. Through evaluation, the effectiveness of the model in real-world trading scenarios can be more reliably assessed.
4 Results
The following formulaic alphas (Table 5) were generated by the LLM. These alphas are grouped by their main focus, including momentum, sentiment, volume, market indices, technical indicators, moving averages, volatility, and various combinations. These alphas incorporate price, volume, and sentiment features to capture diverse trading signals.
Table 6 shows a sample of the thinking process of the DeepSeek model. Unlike other LLMs, DeepSeek generates a thoughtful and detailed reasoning process before presenting results. This approach enhances the explainability of the outcomes, providing investors with clearer insights to inform their financial decisions. For example, when asked to create trading strategies, DeepSeek first analyzes the given features, such as stock prices, technical indicators, and sentiment scores, and identifies their roles in market behavior. It then thoughtfully designs each formula to capture different aspects like momentum, trend strength, volatility, and volume flow, making sure the strategies are diverse and aligned with typical trading logic. Only after carefully laying out this reasoning does DeepSeek present the final set of strategies, allowing users to clearly understand the rationale behind each output.
Table 5: LLM-generated Formulaic Alphas
# Momentum-based formulas alpha1_t = (C_t - O_t) / O_t + 0.5 * Momentum_3 alpha2_t = Momentum_10 * (C_t - SMA_5) alpha3_t = (Momentum_3 + Momentum_10) / 2 alpha4_t = (C_t - SMA_20) * Momentum_3 alpha5_t = Momentum_3 * Momentum_10 # Sentiment-based formulas alpha6_t = S_t * (C_t - O_t) / O_t alpha7_t = Apple_polarity * (C_t - SMA_5) alpha8_t = HSBC_polarity * (C_t - SMA_20) alpha9_t = Pepsi_polarity * Momentum_3 alpha10_t = Tencent_polarity * Momentum_10 # Volume-based formulas alpha11_t = V_t / SMA_20 alpha12_t = OBV * (C_t - O_t) / O_t alpha13_t = V_t * Momentum_3 alpha14_t = V_t * Momentum_10 alpha15_t = OBV * S_t # Index-based formulas alpha16_t = (C_t / Close_Nikkei225) * S_t alpha17_t = (C_t / Close_SP500) * Momentum_3 alpha18_t = (C_t / Close_HSI) * Momentum_10 alpha19_t = Close_Nikkei225 * Close_SP500 * Close_HSI alpha20_t = (Close_Nikkei225 + Close_SP500 + Close_HSI) / 3 # Technical indicator-based formulas alpha21_t = MACD * MACD_Signal alpha22_t = (MACD - MACD_Signal) * S_t alpha23_t = RSI_14 * Momentum_3 alpha24_t = RSI_14 * Momentum_10 alpha25_t = BB_Upper - BB_Lower # Moving average-based formulas alpha26_t = (SMA_5 - SMA_20) * S_t alpha27_t = (EMA_10 - SMA_20) * Momentum_3 alpha28_t = SMA_5 * SMA_20 alpha29_t = EMA_10 * SMA_20 alpha30_t = (SMA_5 + SMA_20) / 2 # Combination formulas alpha31_t = (C_t - SMA_5) / SMA_5 + 0.5 * S_t alpha32_t = (C_t - SMA_20) / SMA_20 + 0.5 * Momentum_3 alpha33_t = (C_t - EMA_10) / EMA_10 + 0.5 * Momentum_10 alpha34_t = (C_t - SMA_5) / SMA_5 + (C_t - SMA_20) / SMA_20 alpha35_t = (C_t - SMA_5) / SMA_5 + (C_t - EMA_10) / EMA_10 # Volatility-based formulas alpha36_t = (High_t - Low_t) / Close_t * S_t alpha37_t = (High_t - Low_t) / Close_t * Momentum_3 alpha38_t = (High_t - Low_t) / Close_t * Momentum_10 alpha39_t = BB_Upper - BB_Lower alpha40_t = (BB_Upper - C_t) / (BB_Upper - BB_Lower) # More combination formulas alpha41_t = (C_t - SMA_5) / SMA_5 + (C_t - SMA_20) / SMA_20 + S_t alpha42_t = (C_t - SMA_5) / SMA_5 + (C_t - EMA_10) / EMA_10 + Momentum_3 alpha43_t = (C_t - SMA_5) / SMA_5 + (C_t - EMA_10) / EMA_10 + Momentum_10 alpha44_t = (C_t - SMA_5) / SMA_5 + (C_t - SMA_20) / SMA_20 + Momentum_3 alpha45_t = (C_t - SMA_5) / SMA_5 + (C_t - SMA_20) / SMA_20 + Momentum_10 # Advanced combination formulas alpha46_t = (C_t - SMA_5) / SMA_5 + (C_t - SMA_20) / SMA_20 + S_t + Momentum_3 alpha47_t = (C_t - SMA_5) / SMA_5 + (C_t - EMA_10) / EMA_10 + S_t + Momentum_3 alpha48_t = (C_t - SMA_5) / SMA_5 + (C_t - EMA_10) / EMA_10 + S_t + Momentum_10 alpha49_t = (C_t - SMA_5) / SMA_5 + (C_t - SMA_20) / SMA_20 + S_t + Momentum_10 alpha50_t = (C_t - SMA_5) / SMA_5 + (C_t - SMA_20) / SMA_20 + S_t + Momentum_3 + Momentum_10
(a) Apple
(b) HSBC
(c) Tencent
(d) Toyota
(e) Pepsi
Figure 4: Mutual Information (in Blue) and Information Coefficient (in Red) of LLM-generated Alphas
(a) Apple
(b) HSBC
(c) Tencent
(d) Toyota
(e) Pepsi
Figure 5: Feature Importance of LLM-generated Alphas
Table 6: Deepseek Thinking Process
Alright, the user wants me to generate 50 alpha formulas using the given stock features. They specified that the formulas should be mathematical expressions combining these features and returned in Python syntax with specific variable names. First, I need to understand what each feature represents. Features like Open, High, Low, Close, Volume are standard OHLCV data. Then there are indices like Nikkei 225, S&P 500, and HSI, which are market indicators. Sentiment and polarities for specific companies (Apple, HSBC, Pepsi, Tencent) are sentiment analysis metrics. The rest are technical indicators: SMA, EMA, Momentum, RSI, MACD, and OBV. I should start by considering different categories of formulas. Maybe some based on momentum, others on volatility, sentiment, volume, etc. This way, the formulas will cover a broad range of strategies. For momentum-based formulas, I can use the Momentum indicators, maybe combine them with price changes. For example, (C_t - O_t) / O_t multiplied by Momentum_3. That makes sense because it ties the day’s price movement to recent momentum. Next, I can think about volatility. Using High and Low prices, I can calculate daily ranges. Maybe (High_t - Low_t) / Close_t to get a volatility measure. Then combine this with moving averages or sentiment. Sentiment is another key area. The user provided overall sentiment and specific company polarities. I should incorporate these to capture market mood. For example, using S_t (sentiment) multiplied by some function of price or volume. Volume is important for confirmation. Formulas that include V_t can indicate strength behind price movements. Maybe V_t compared to a moving average of volume, or V_t in relation to price changes. I should also consider combinations of different indicators. For instance, using MACD and its signal line to generate a crossover signal, then scaling it by sentiment or volume. Technical indicators like RSI can signal overbought or oversold conditions. I can use RSI levels to create mean-reversion strategies, perhaps combined with other indicators for confirmation. I need to ensure each formula is unique and covers different aspects. Maybe some formulas focus on single indicators, while others combine multiple. For example, using BB_Upper and BB_Lower to create a volatility-based strategy, or combining SMA_5 and SMA_20 to look for crossovers. I also need to think about the structure of each formula. They should be mathematical expressions, so I’$l$ use operations like addition, subtraction, multiplication, division, and maybe exponents or logarithms where appropriate. Let me start drafting some formulas. For the first one, a simple momentum strategy: (C_t - O_t) / O_t. That’s the daily return. Then, I can add a sentiment factor, maybe 0.5 * S_t, to incorporate market mood. Another idea: using the relative strength of the stock compared to an index. For example, (C_t / Close_Nikkei225) to see if the stock is outperforming the index. I should also use technical indicators in creative ways. For example, MACD_t - MACD_Signal_t gives the difference between the two lines, which can indicate a bullish or bearish signal. Volume can be used to confirm the strength of a trend. So, V_t / SMA_Volume might show if current volume is higher than average, indicating strong movement. I need to make sure I use all the provided features. That means including each feature in at least a few formulas. For example, using Apple_polarity in some strategies, HSBC_polarity in others, etc. I can also create formulas that normalize the data, like (C_t - SMA_5) / SMA_20, to see how the close price relates to different moving averages. For each formula, I’$l$ assign a variable name like alpha1_t, alpha2_t, etc., and express it in Python syntax using the specified variable abbreviations. I should also vary the complexity. Some formulas can be simple two-term expressions, while others might have three or more terms, possibly with coefficients to weight different factors. I need to ensure that each formula is plausible and could be used in a real trading strategy. They should be based on common trading logics, like trend-following, mean-reversion, momentum, etc. Let me think of some more. Using RSI in a formula, perhaps as 1 / RSI_t to indicate oversold conditions. Or combining RSI with momentum for a dual approach. I should also consider combinations of different moving averages, like EMA_10 and SMA_20, to create crossovers or compare their values. For the sentiment variables, I can use them as direct multipliers or add them after scaling. For example, S_t * (C_t - O_t) / O_t to tie sentiment to price movement. I need to make sure that all 50 formulas are unique and don’t repeat the same logic. Each should offer a different perspective or combination of features. Finally, I’$l$ review each formula to ensure they’re mathematically sound and use the correct variable names as per the user’s instructions.
4.1 MI and IC Analysis
Figure 4 presents the mutual information (MI) and the information coefficient (IC) of the fifty LLM-generated alphas in five companies. In the figure, the blue bars indicate the MI values, while the red bars represent the IC values. Together, they offer insight into both the statistical dependence between the alphas and returns (MI), and the directional predictive power (IC).
Apple shows the strongest overall signal among the five companies, with a larger number of alphas exhibiting positive MI values. For example, Alpha 3, 12, 14, 20, 23, 24, 25, 29, and 31 all have MI values exceeding 0.04, suggesting a moderate to strong dependency between alpha signals and Apple returns. However, their IC values vary and several of them are negative. This indicates that while the alphas are related to future returns, they may not correctly predict the direction. In particular, Alpha 3, 24, 33, 38, 48, 49, and 50 have higher IC values, highlighting a few strong directional signals.
Pepsi also shows relatively high MI values for many alphas, which implies that these features contain some information about future returns. However, most IC values for Pepsi are negative, suggesting that although the signals are related to the returns, their predictive direction is often incorrect or weak.
In contrast, HSBC stands out for its high IC values. Alphas such as Alpha 11, 14, 18, 24, 33, and 38 achieve IC values above 0.04, indicating a strong and consistent predictive direction. This suggests that LLM-generated alphas are more effective in forecasting HSBC return movements compared to other stocks.
On the other hand, Tencent exhibits generally low MI values and mostly negative IC values. This implies that LLM-generated alphas show a weak dependence on Tencent’s stock returns and also fail to capture directional accuracy.
Toyota also shows a few alphas with notably strong predictive signals. In particular, Alpha 2, 4, and 5 achieve IC values exceeding 0.06, indicating a strong and reliable correlation with future stock returns. These high IC values suggest that these alphas consistently capture the correct direction of Toyota’s price movements. However, the majority of the alphas for Toyota fall within the IC range of 0.01 to 0.03, which reflects relatively weak but potentially usable signals. This distribution implies that while only a few alphas offer strong directional accuracy, many still exhibit a mild level of predictive power.
In general, these results suggest that the predictive quality of LLM-generated alphas can vary significantly between companies and that high mutual information does not always translate into high directional predictive power.
4.2 Feature Importance
Figure 5 displays the LightGBM feature importance scores for each LLM-generated alpha across the five selected companies. The results indicate that Alpha 25 stands out as the most influential feature for Apple, Toyota, and Pepsi, suggesting that this alpha consistently contributes valuable information to the prediction models for these companies. In contrast, Alpha 21 holds the highest importance for HSBC and Tencent, implying that different alphas capture distinct patterns depending on the company. On the other hand, Alpha 39 appears to have minimal importance across all five companies, indicating limited predictive relevance in most cases. While there are differences in which alphas are prioritized, the overall distribution of importance scores does not show extreme variation, suggesting a relatively balanced reliance on multiple alphas rather than dominance by a few. This diversity highlights that the predictive power of each alpha may depend heavily on the underlying characteristics of each stock.
4.3 Comparison of PPO-Adjusted Strategy and Market Benchmark
Table 7 presents the performance metrics of both the PPO-adjusted strategy and the market benchmark. For each stock, the framework is executed ten times using non-deterministic inference to allow stochastic behavior, which is an inherent feature of the PPO adjustment and important for capturing the variability in trading outcomes. The results are reported as the mean and standard deviation in the table. Reporting both statistics provides a measure of the average performance as well as the variability across runs, thus reflecting the stability and robustness of the framework.
The PPO-adjusted strategy shows strong performance across most companies compared to their respective market benchmarks. Apple, HSBC, Pepsi, and Tencent all achieved significantly higher cumulative returns, suggesting that the strategy was effective in capturing upside potential. Toyota, however, experienced a relatively lower return, underperforming both the other companies and its benchmark. This may indicate that the strategy struggled in the Japanese market or that market conditions during the test period were less favorable.
When looking at risk-adjusted performance through the Sharpe ratio, the PPO-adjusted strategy maintained consistently high values for Apple, Pepsi, and Tencent. This implies not only strong returns but also relatively low volatility in these cases. On the other hand, Toyota had the lowest Sharpe ratio, reinforcing the poor performance observed in its cumulative return. Drawdown values were low for most companies, indicating limited downside risk. However, Toyota again stood out with a larger drawdown, which aligns with its lower return and Sharpe ratio. Overall, the PPO strategy appears to be effective in various markets, though its success may vary depending on the specific company or market environment.
The standard deviations provide insight into the consistency of the results. For Apple and Pepsi, the standard deviations are very small relative to the mean, indicating highly consistent performance. HSBC and Tencent exhibit moderate variability, with fluctuations larger than those of the U.S. stocks. In contrast, Toyota shows relatively large standard deviations compared to its mean values, suggesting that its results are less stable. This instability may be due to several factors: the LLM-generated alphas may capture weaker or noisier signals for Toyota, making the PPO-adjusted weights fluctuate more; Toyota trades in the Nikkei 225, which can have higher volatility and lower liquidity compared to the U.S. markets.
4.4 Comparison of PPO-Adjusted Strategy and Equal-Weighted Strategy
Table 8 shows the results of an equal-weighted strategy constructed using the same set of alphas. In general, the PPO-adjusted strategy demonstrates superior performance across most stocks except Toyota when compared to the equal-weighted strategy. Specifically, for most of the time, the PPO-adjusted approach achieves higher cumulative returns and Sharpe ratios, indicating both stronger profitability and more efficient risk-adjusted performance. Furthermore, the PPO-adjusted strategy exhibits lower maximum drawdowns, reflecting improved downside protection.
In contrast, the equal-weighted strategy produces mixed results, with lower returns and higher drawdowns for most stocks, highlighting the limitations of assigning equal weight to all alphas without considering their predictive contribution and market dynamics. These comparisons suggest that the PPO-adjusted framework effectively identifies and emphasizes more informative alphas, leading to a more robust and stable trading performance relative to the equal-weighted approach.
Table 7: PPO-Adjusted Strategy vs. Market Benchmark (with 50 Alphas), Mean (Standard Deviation)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return 1.6817 (0.0619) 0.4657 (0.0554) 0.6272 (0.0331) 0.6245 (0.0632) 0.0299 (0.0398)
Sharpe Ratio 1.9998 (0.0169) 0.8131 (0.0691) 1.4319 (0.0434) 1.1440 (0.1052) 0.1011 (0.0786)
Drawdown -0.0101 (0.0007) -0.2259 (0.0099) -0.0067 (0.0001) -0.0810 (0.0208) -0.2801 (0.0076)
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
Table 8: Equal-Weighted Strategy (with 50 Alphas)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return -0.3200 -0.9059 -0.1726 -0.7074 0.1026
Sharpe Ratio -0.3601 -0.3580 -0.2879 -0.2032 0.1586
Drawdown -0.3663 -0.9489 -0.2922 -0.8741 -0.3571
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
Table 9: PPO-Adjusted Strategy Using Low-Correlation Alphas vs. Market Benchmark, Mean (Standard Deviation)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return 1.6641 (0.1755) 0.4983 (0.0660) 0.6319 (0.0775) 0.5956 (0.0722) 0.0842 (0.0579)
Sharpe Ratio 2.0002 (0.0692) 0.8388 (0.0776) 1.4259 (0.0622) 1.1191 (0.1018) 0.2072 (0.1083)
Drawdown -0.0113 (0.0016) -0.2210 (0.0121) -0.0066 (0.0002) -0.0810 (0.0183) -0.2344 (0.0288)
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
Table 10: PPO-Adjusted Strategy Using High-Contribution Alphas vs. Market Benchmark, Mean (Standard Deviation)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return 1.6072 (0.1093) 0.5199 (0.0991) 0.6023 (0.0441) 0.5634 (0.1200) 0.0741 (0.0395)
Sharpe Ratio 1.9890 (0.0626) 0.8791 (0.1215) 1.3996 (0.0296) 1.0620 (0.1735) 0.1894 (0.0765)
Drawdown -0.0113 (0.0016) -0.2286 (0.0253) -0.0076 (0.0022) -0.0919 (0.0263) -0.2625 (0.0217)
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
Table 11: PPO-Adjusted Strategy Using 30 Randomly Selected Alphas vs. Market Benchmark, Mean (Standard Deviation)
Metric Apple HSBC Pepsi Tencent Toyota
Cum Return 1.7276 (0.0647) 0.4853 (0.0451) 0.6176 (0.0225) 0.6041 (0.0755) 0.0410 (0.0291)
Sharpe Ratio 2.0133 (0.0255) 0.8399 (0.0534) 1.4271 (0.0367) 1.1129 (0.1132) 0.1238 (0.0565)
Drawdown -0.0109 (0.0009) -0.2248 (0.0061) -0.0068 (0.0006) -0.0937 (0.0307) -0.2664 (0.0137)
Market Return 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.3476 (S&P 500) 0.0033 (Hang Seng) 0.4081 (Nikkei 225)
5 Further Discussion
5.1 Impact of Different Alpha Selection Settings on Trading Performance
Table 12: Number of Alphas Selected by Different Reduction Methods for Each Company
Method Apple HSBC Pepsi Tencent Toyota
Low-Correlation Selection 13 14 14 18 10
High-Contribution Selection 10 10 10 10 10
Random Selection 30 30 30 30 30
To evaluate how the number of formulaic alphas influences trading performance, we conduct a more detailed analysis by adjusting the number of LLM-generated alphas fed into the reinforcement learning module. Three different methods are used. For the first approach, we reduce redundancy by measuring the pairwise correlations between them. Highly correlated alphas often carry overlapping information, which may not contribute much to the learning process and could even introduce noise. To address this, we remove alphas with the absolute value of pairwise correlation coefficients (Figure 6) exceeding a threshold of 0.7 (Figure 7), a common choice in quantitative research. This helps retain only the most diverse and informative signals for the learning process. For the second approach, we choose the top ten alphas with the highest contribution in terms of the LightGBM feature importance (Gain) (Figure 5). For the third approach, thirty alphas are chosen randomly for each company.
Table 9 presents the results when using low-correlation alphas. Comparing the PPO strategy using all 50 alphas with the version that applies low-correlation filtering, we observe that reducing redundancy among alphas affects stocks differently. For Apple and Tencent, performance declines slightly in terms of returns or risk-adjusted metrics, suggesting that some removed alphas may have contributed useful information. In contrast, HSBC, Pepsi, and Toyota show modest improvements, indicating that the elimination of highly correlated alphas can slightly enhance the signal quality for these stocks. In general, the low-correlation approach maintains the ability of the strategy to outperform market benchmarks while slightly increasing variability for certain stocks, reflecting a trade-off between signal diversity and stability.
Table 10 shows the performance using high-contribution alphas. Specifically, Apple, Pepsi, and Tencent experience slight declines in cumulative returns and Sharpe ratios, along with higher drawdowns, suggesting that focusing solely on the top-contributing alphas may reduce the diversity of useful signals for these stocks. In contrast, HSBC and Toyota see improvements in both return and risk-adjusted performance, indicating that high-contribution alphas provide particularly strong signals for these assets. Overall, while the high-contribution approach maintains the ability to outperform market benchmarks, it does not consistently improve performance across all stocks and may sacrifice stability for certain assets.
Table 11 presents the performance of the PPO-adjusted strategy when using thirty randomly selected alphas. Apple continues to exhibit strong performance, achieving the highest cumulative return and Sharpe ratio among all five stocks. In contrast, Toyota still performs the worst in this setting. Notably, HSBC shows improved performance compared to the scenario using all fifty alphas in terms of return and Sharpe ratio. Overall, despite the reduced number of alphas, all stocks still outperform their respective market benchmarks.
In general, although the impact varies across individual stocks, the different alpha selection settings do not substantially alter the overall results.
(a) Apple
(b) HSBC
(c) Tencent
(d) Toyota
(e) Pepsi
Figure 6: Correlation Heatmaps of 50 LLM-generated Alphas
(a) Apple
(b) HSBC
(c) Tencent
(d) Toyota
(e) Pepsi
Figure 7: Correlation Heatmaps of LLM-generated Alphas ($\left|\text{Corr}\right|$ $\leq 0.7$)
5.2 Impact of Prompt Information on Trading Performance
In this subsection, we examine how formulaic alphas generated with different prompt information influence the performance of the proposed RL strategy. Originally, the prompt is fed a pandas DataFrame in JSON format containing various features of all companies. Here, we consider two additional scenarios: (1) providing the prompt with only feature names and no additional data, instructing the model to generate alphas based solely on its own knowledge; and (2) providing the prompt with partial information, such as data from a single company.
For simplicity, the analysis focuses solely on Apple. In the partial information scenario, only data from Apple is provided. Table 13 demonstrates the PPO strategy performance under different scenarios. Based on the results, we observe that across all scenarios, formulaic alphas generated with varying prompt information produce strong performance under the proposed RL framework, with cumulative returns exceeding the market benchmark and Sharpe ratios above one, demonstrating the robustness and effectiveness of the framework.
Table 13: Performance of PPO-Adjusted Strategy on Apple under Different Prompt Information. Mean (Standard Deviation).
Metric Full Information Partial Information Feature Name Only
Cumulative Return 1.6817 (0.0619) 1.4312 (0.1342) 1.2134 (0.0848)
Sharpe Ratio 1.9998 (0.0169) 1.9715 (0.0622) 1.9540 (0.0360)
Drawdown -0.0101 (0.0007) -0.0130 (0.0011) -0.0116 (0.0012)
Market Return 0.3476 (S&P 500) 0.3476 (S&P 500) 0.3476 (S&P 500)
5.3 Impact of Sentiment on Trading Performance
To evaluate the impact of sentiment on trading performance, we consider three settings: (1) sentiment of the target company only, (2) sentiment of all companies, and (3) no sentiment. For each target company, fifty unique alphas are generated in each scenario. As demonstrated in Section 5.2, using feature names alone is sufficient to achieve strong trading performance. Therefore, in this section, we simplify the prompts by including only feature names.
Table 14 to 18 report the performance of the proposed PPO strategy for all five companies under different sentiment settings. The results suggest that incorporating sentiment features within the PPO framework does not lead to substantial differences in trading performance. Across all three sentiment settings for each target company, the proposed strategy consistently delivers strong results.
Table 14: Performance of PPO-Adjusted Strategy on Apple under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 1.2499 (0.8023) 1.4797 ( 0.1156) 1.4459 (0.1098)
Sharpe Ratio 1.9676 (0.0344) 2.0134 (0.0421) 1.9933 (0.0387)
Drawdown -0.0133 (0.0013) -0.0084 (0.0019) -0.0110 (0.0020)
Market Return 0.3476 (S&P 500) 0.3476 (S&P 500) 0.3476 (S&P 500)
Table 15: Performance of PPO-Adjusted Strategy on HSBC under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 0.6093 (0.0455) 0.4460 (0.0525) 0.5955 (0.0481)
Sharpe Ratio 1.2690 (0.2100) 0.9343 (0.1959) 1.1638 (0.2021)
Drawdown -0.1014 (0.0239) -0.1597 (0.0311) -0.1489 (0.0283)
Market Return 0.0033 (Hang Seng) 0.0033 (Hang Seng) 0.0033 (Hang Seng)
Table 16: Performance of PPO-Adjusted Strategy on Pepsi under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 0.4124 (0.0396) 0.7643 (0.0514) 0.7644 (0.0437)
Sharpe Ratio 1.7663 (0.1956) 1.7786 (0.1822) 1.7972 (0.1891)
Drawdown -0.0454 (0.0111) -0.0051 (0.0070) -0.0047 (0.0064)
Market Return 0.3476 (S&P 500) 0.3476 (S&P 500) 0.3476 (S&P 500)
Table 17: Performance of PPO-Adjusted Strategy on Tencent under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 0.6136 (0.0486) 0.7163 (0.0436) 0.7067 (0.0513)
Sharpe Ratio 1.1961 (0.1821) 1.3742 (0.1953) 1.2433 (0.1779)
Drawdown -0.0515 (0.0120) -0.0504 (0.0132) -0.0484 (0.0112)
Market Return 0.0033 (Hang Seng) 0.0033 (Hang Seng) 0.0033 (Hang Seng)
Table 18: Performance of PPO-Adjusted Strategy on Toyota under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 0.0272 (0.0554) 0.1088 (0.0622) 0.0102 (0.0512)
Sharpe Ratio 0.0982 (0.0326) 0.2549 (0.0451) 0.0641 (0.0287)
Drawdown -0.2735 (0.0413) -0.2203 (0.0370) -0.2651 (0.0395)
Market Return 0.4081 (Nikkei 225) 0.4081 (Nikkei 225) 0.4081 (Nikkei 225)
6 Conclusion
This research explores the use of LLMs for generating formulaic alphas and applies PPO to optimize their weights for trading decisions. Experimental results show that the PPO-adjusted strategy achieves significantly higher cumulative returns than traditional benchmarks such as the S&P 500, Hang Seng, and Nikkei 225 for most target stocks. In addition to higher returns, the strategy also demonstrates strong risk-adjusted performance, as reflected in consistently high Sharpe ratios across companies like Apple and Pepsi. Moreover, the strategy maintains relatively low drawdown levels, indicating better downside protection compared to market indices. Additionally, the PPO strategy outperforms an equal-weighted alpha portfolio for most of the stocks.
Reducing the number of alphas leads to slight improvements in both cumulative return and Sharpe ratio for certain stocks. However, for others, their performance may worsen after reducing the number of alphas. This indicates that selecting a well-balanced set of informative and diverse alphas can enhance both stability and profitability. The results underscore the potential of combining LLMs with reinforcement learning to create effective and interpretable trading strategies. Future work will explore alternative LLM architectures, additional reinforcement learning methods, and broader applications across different markets and asset classes.